OpenStudy (anonymous):

what is 2(y-6x) = 12 ??

OpenStudy (marigirl):

what subject are you studying in maths? if it is graphing, we can express the equation in terms of y 2(y-6x)=12 y-6x=6 (divide both sides by 2) y=6x+6 and we can graph that....

OpenStudy (anonymous):

Thank you! I'm studying algebra 1 and I'm supposed to convert it so it will match y=mx+b

OpenStudy (anonymous):

But what if it say y-x=5 ?

OpenStudy (skullpatrol):

add x to both sides

OpenStudy (anonymous):

and that would be 5X?

OpenStudy (skullpatrol):

what do you mean "5X"?

OpenStudy (skullpatrol):